Circular No. 8703

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 GZ_{2} (SPACEWATCH) An apparently asteroidal object discovered by Spacewatch (discovery

observation below) was reported already on Apr. 7.9 UT to be slightly diffuse in follow-up images obtained in strong moonlight by J. Tichá and M. Tichý with the 1.06-m telescope at Kleť. With no other reports on physical appearance available, the object was announced as 2006 GZ_{2} on Apr. 10 (MPEC 2006-G38; also MPS 168881), but on the following day, J. Montani reported that his inspection of the Spacewatch images from Apr. 7.2, 8.2, and 9.2 all show the object to have a coma of diameter 5′′–6′′ and no tail. R. S. McMillan adds that Spacewatch images taken by T. H. Bressi on Apr. 18.4 also show the object to be diffuse with a 4′′–5′′ coma.

2006 UT  α_{2000}  δ_{2000}  Mag.
Apr. 7.18152  12^{h}53^{m}41.^{s}82  −0^{o}16′27′′.9  20.1

More recent astrometry, the following parabolic orbital elements, and an ephemeris appear on MPEC 2006-H10.

T = 2006 Aug. 22.050 TT  ω = 191.^{o}508
Ω = 355.320 }2000.0
q = 3.30214 AU  i = 168.683

COMET 73P/SCHWASSMANN-WACHMANN Futher to IAUC 8701, numerous additional components to 73P have

been observed on three or more nights and thus given designations (‘T’–‘Z’, and ‘AA’–‘AM’, although component ‘Y’ has only two nights of astrometry so far) and first announced on MPECs 2006-G10 and 2006-H03.

COMET P/2006 G1 (McNAUGHT) Additional observations published on MPEC 2006-G43 and MPC 56610

show that this comet (cf. IAUC 8699) is indeed of short period, with the following improved orbital elements:

T = 2006 Aug. 20.7210 TT  ω = 314.^{o}8793
e = 0.455391  Ω = 299.0576 }2000.0
q = 2.621261 AU  i = 18.5516
a = 4.813108 AU n^{o} = 0.0933395 P = 10.56 years
2006 April 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green